# Landau.bib

### v1

@@ 3.2.2 Approximated chain rule
    $\lim_{\beta\to 0}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}|\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}|^{2}{\,\rm d}\eta{\,\rm d}t$ (3.41)
    $=\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}\frac{|\nabla_{v}f^{\alpha}|^{2}}{f^{\alpha}}{\,\rm d}\eta{\,\rm d}t.$
@@
    We first show the convergence (3.40). By using of the identity $(v-w)\times\nabla M_{\beta}(v-w)=0$, we have
    $v\times\big{(}\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}\big{)}=v\times\int_{\mathbb{R}^{3}}\nabla_{v}M_{\beta}(v-w)\log f^{\alpha,\beta}(w){\,\rm d}w$
    $=\int_{\mathbb{R}^{3}}w\times\nabla_{v}M_{\beta}(v-w)\log f^{\alpha,\beta}(w){\,\rm d}w=\big{(}v\times\nabla_{v}\log f^{\alpha,\beta}\big{)}*_{v}M_{\beta}.$
@@
    $|\big{(}v\times\nabla_{v}\log f^{\alpha,\beta}\big{)}*_{v}M_{\beta}|^{2}$ (3.42)
[…]
    ×$\quad\times\int_{\mathbb{R}^{3}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\big{|}w\times\nabla_{w}\log f^{\alpha,\beta}\big{|}^{2}{\,\rm d}w$
[…]
@@
    $\int_{0}^{T}\int_{\mathbb{R}^{3\times 4}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v\times(\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta})|^{2}{\,\rm d}\eta{\,\rm d}t$
[…]
    ×$\quad\times\Big{(}\int_{\mathbb{R}^{3}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\Big{|}\frac{w\times\nabla_{w}f^{\alpha,\beta}}{f^{\alpha,\beta}}\Big{|}^{2}{\,\rm d}w\Big{)}{\,\rm d}x{\,\rm d}v{\,\rm d}t$
[…]
@@
    $\langle v\rangle^{\gamma}\frac{|v\times\nabla_{v}f^{\alpha}*_{v}M_{\beta}|^{2}}{f^{\alpha}*_{v}M_{\beta}}\leq\langle v\rangle^{\gamma}M_{\beta}*_{v}M_{\alpha}*_{x}\frac{|v\times\nabla_{v}f|^{2}}{f}.$ (3.43)
@@
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\langle v\rangle^{\gamma}M_{\beta}*_{v}M_{\alpha}*_{x}\frac{|v\times\nabla_{v}f|^{2}}{f}{\,\rm d}x{\,\rm d}v\lesssim\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\langle v\rangle^{\gamma}\frac{|v\times\nabla_{v}f|^{2}}{f}{\,\rm d}x{\,\rm d}v,$ (3.44)
@@
    $\int_{0}^{T}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big{)}$
    ×$\quad\times f^{\alpha}\langle v\rangle^{-\gamma}\big{(}M_{\beta}*_{v}(\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2})\big{)}{\,\rm d}x{\,\rm d}v{\,\rm d}t$
[…]
@@
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f^{\alpha,\beta}|^{2}}{f^{\alpha,\beta}}{\,\rm d}x{\,\rm d}v\lesssim\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f|^{2}}{f}{\,\rm d}x{\,\rm d}v.$
@@
    $\Big{|}\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}(\log f^{\alpha})\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r\Big{|}\leq\frac{1}{2}\int_{0}^{T}\mathcal{D}(f^{\alpha}){\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f^{\alpha},U^{\alpha}){\,\rm d}t.$
@@
    $\lim_{\delta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $=(1-\theta)\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{3}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r.$
@@
    $\leq C(\theta,\alpha)\Big{(}\int_{\mathbb{R}^{3}}\big{(}\langle v\rangle+\langle x\rangle\big{)}|u|M_{\beta}(u)M_{\alpha}(y)f^{\delta}(x-y,v-u){\,\rm d}y\Big{)}$
@@
    $\Big{|}\int_{s}^{t}\int_{\mathbb{R}^{12}}\big{(}\langle v\rangle+\langle x\rangle\big{)}|u|M_{\beta}(u)M_{\alpha}(y)|f^{\delta}-f|(x-y,v-u){\,\rm d}y{\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big{|}$
    $\leq\|\langle v\rangle^{2}M_{\beta}\|_{L^{1}}\|\langle x\rangle^{2}M_{\alpha}\|_{L^{1}}\Big{|}\int_{0}^{T}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\big{(}\langle v\rangle+\langle x\rangle\big{)}(f^{\delta}_{t}-f_{t}){\,\rm d}x{\,\rm d}v{\,\rm d}t\Big{|},$
    which vanishes as $\delta\to 0$, since $f\in L^{1}([0,T];L^{1}_{1,1}(\mathbb{R}^{3}\times\mathbb{R}^{3}))$. By the dominated convergence theorem, we obtain the limit as $\delta\to 0$.
@@
    $\Big{|}\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{3}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big{|}$
    $\leq C(\alpha,\beta)\int_{s}^{t}\int_{\mathbb{R}^{12}}\big{(}\langle v\rangle+\langle x\rangle+\mathcal{E}_{2,0}(f_{t})^{\frac{1}{2}}\big{)}$
    ×$\quad\times|u|M_{\beta}(u)M_{\alpha}(y)f(x-y,v-u){\,\rm d}y{\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
@@
    $\lim_{\beta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{3}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r=0.$
@@
    $\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\log f^{\alpha,\beta}_{r}(v)\int_{\mathbb{R}^{3}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $=-\int_{s}^{t}\int_{\mathbb{R}^{3\times 3}}\frac{\nabla_{x}f^{\alpha,\beta}_{r}(v)}{f^{\alpha,\beta}_{r}(v)}\cdot uf^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r.$
@@
    $\Big{|}\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\nabla_{x}f^{\alpha,\beta}_{r}(v)}{f^{\alpha,\beta}_{r}(v)}\cdot\int_{\mathbb{R}^{3}}uf^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big{|}$
    $\leq\Big{(}\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{|\nabla_{x}f^{\alpha,\beta}_{r}|^{2}}{f^{\alpha,\beta}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big{)}^{\frac{1}{2}}\times$
    ×$\quad\times\Big{(}\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{(\int_{\mathbb{R}^{3}}|u|f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u)^{2}}{\int_{\mathbb{R}^{3}}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u}{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big{)}^{\frac{1}{2}}.$
@@
    $\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{(\int_{\mathbb{R}^{3}}|u|f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u)^{2}}{\int_{\mathbb{R}^{3}}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
[…]
@@
    $\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{|\nabla_{x}f^{\alpha,\beta}_{r}|^{2}}{f^{\alpha,\beta}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r$ $\leq\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{|\nabla_{x}f^{\alpha}_{r}|^{2}}{f^{\alpha}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
@@
    $|\int_{s}^{t}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\log f^{\alpha,\beta}_{r}(v)\int_{\mathbb{R}^{3}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r|$
@@
[…]
@@
We left to show the $\alpha,\,\beta\to 0$ limits. Since $\mathcal{H}(f^{\alpha,\beta}_{t})$ is lower semicontinuous and $f^{\alpha,\beta}_{t}\to f_{t}$ in $L^{1}_{2,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ as $\alpha,\,\beta\to 0$, we have
@@
$\mathcal{H}\Big{(}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f_{t}(x-y,v-u)M_{\alpha}(y)M_{\beta}(u){\,\rm d}y{\,\rm d}u\Big{)}\leq\mathcal{H}(f_{t}).$

### v2

@@ 3.2.2 Approximated chain rule
    $\lim_{\beta\to 0}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}|\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}|^{2}{\,\rm d}\eta{\,\rm d}t$ (3.41)
    $=\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}\frac{|\nabla_{v}f^{\alpha}|^{2}}{f^{\alpha}}{\,\rm d}\eta{\,\rm d}t.$
@@
    We first show the convergence (3.40). By using of the identity $v-w,\nabla M_{\beta}(v-w)]_{ij}=0$, we have
    $[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}=\Big[v,\int_{\mathbb{R}^{d}}\nabla_{v}M_{\beta}(v-w)\log f^{\alpha,\beta}(w){\,\rm d}w\Big]$
    $=\int_{\mathbb{R}^{d}}[w,\nabla_{v}M_{\beta}(v-w)\log f^{\alpha,\beta}(w)]_{ij}{\,\rm d}w=\big([v,\nabla_{v}\log f^{\alpha,\beta}]_{ij}\big)*_{v}M_{\beta}.$
@@
    $|\big([v,\nabla_{v}\log f^{\alpha,\beta}]_{ij}\big)*_{v}M_{\beta}|^{2}$ (3.42)
    $\leq{}$ $\Big(\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{-\gamma}{\,\rm d}w\Big)$
    $\quad\times\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\big|[w,\nabla_{w}\log f^{\alpha,\beta}]_{ij}\big|^{2}{\,\rm d}w$
    $\lesssim{}$ $\langle v\rangle^{-\gamma}\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\Big|\frac{[w,\nabla_{w}f^{\alpha,\beta}]_{ij}}{f^{\alpha,\beta}}\Big|^{2}{\,\rm d}w,$
@@
    $\int_{0}^{T}\int_{\mathbb{R}^{3\times 4}}\kappa f^{\alpha}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|[v,\nabla_{v}\log f^{\alpha,\beta}*_{v}M_{\beta}]_{ij}|^{2}{\,\rm d}\eta{\,\rm d}t$
    $\lesssim{}$ $\int_{0}^{T}\int_{\mathbb{R}^{2d}}\Big(\int_{\mathbb{R}^{2d}}f^{\alpha}_{*}|v-v_{*}|^{\gamma}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big)f^{\alpha}\langle v\rangle^{-\gamma}$
    $\quad\times\Big(\int_{\mathbb{R}^{d}}M_{\beta}(v-w)\langle w\rangle^{\gamma}\Big|\frac{[w,\nabla_{w}f^{\alpha,\beta}]_{ij}}{f^{\alpha,\beta}}\Big|^{2}{\,\rm d}w\Big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\lesssim{}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,0}(f_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}f^{\alpha}M_{\beta}*_{v}\Big(\langle v\rangle^{\gamma}\times\Big|\frac{[v,\nabla_{v}f^{\alpha,\beta}]_{ij}}{f^{\alpha,\beta}}\Big|^{2}\Big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $={}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,0}(f_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\times\frac{|[v,\nabla_{v}f^{\alpha,\beta}]_{ij}|^{2}}{f^{\alpha,\beta}}{\,\rm d}x{\,\rm d}v{\,\rm d}t,$
@@
    $\langle v\rangle^{\gamma}\frac{|[v,\nabla_{v}f^{\alpha}*_{v}M_{\beta}|^{2}]_{ij}}{f^{\alpha}*_{v}M_{\beta}}\leq\langle v\rangle^{\gamma}M_{\beta}*_{v}M_{\alpha}*_{x}\frac{|[v,\nabla_{v}f]_{ij}|^{2}}{f}.$ (3.43)
@@
    $\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}M_{\beta}*_{v}M_{\alpha}*_{x}\frac{|[v,\nabla_{v}f]_{ij}|^{2}}{f}{\,\rm d}x{\,\rm d}v\lesssim\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|[v,\nabla_{v}f]_{ij}|^{2}}{f}{\,\rm d}x{\,\rm d}v,$ (3.44)
@@
    $\int_{0}^{T}\int_{\mathbb{R}^{2d}}\Big(\int_{\mathbb{R}^{2d}}f^{\alpha}_{*}|v-v_{*}|^{\gamma}|v_{*}|^{2}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big)$ (3.45)
    $\quad\times f^{\alpha}\langle v\rangle^{-\gamma}\big(M_{\beta}*_{v}(\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2})\big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\leq{}$ $\int_{0}^{T}C^{\pm}_{\gamma,2}(f^{\alpha}_{t})\int_{\mathbb{R}^{2d}}f^{\alpha}M_{\beta}*_{v}\big(\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2}\big){\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\leq{}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,2}(f^{\alpha}_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta}\langle v\rangle^{\gamma}|\nabla_{v}\log f^{\alpha,\beta}|^{2}{\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $\leq{}$ $\sup_{t\in[0,T]}C^{\pm}_{\gamma,2}(f_{t})\int_{0}^{T}\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f^{\alpha,\beta}|^{2}}{f^{\alpha,\beta}}{\,\rm d}x{\,\rm d}v{\,\rm d}t,$
@@
    $\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f^{\alpha,\beta}|^{2}}{f^{\alpha,\beta}}{\,\rm d}x{\,\rm d}v\lesssim\int_{\mathbb{R}^{2d}}\langle v\rangle^{\gamma}\frac{|\nabla_{v}f|^{2}}{f}{\,\rm d}x{\,\rm d}v.$
@@
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}(\log f^{\alpha})\cdot U^{\alpha}{\,\rm d}\eta{\,\rm d}r\Big|\leq\frac{1}{2}\int_{0}^{T}\mathcal{D}(f^{\alpha}){\,\rm d}t+\frac{1}{2}\int_{0}^{T}\mathcal{A}(f^{\alpha},U^{\alpha}){\,\rm d}t.$
@@
    $\lim_{\delta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{2d}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $=(1-\theta)\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r.$
@@
    $\leq C(\theta,\alpha)\Big(\int_{\mathbb{R}^{d}}\big(\langle v\rangle+\langle x\rangle\big)|u|M_{\beta}(u)M_{\alpha}(y)f^{\delta}(x-y,v-u){\,\rm d}y\Big)$
@@
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{4d}}\big(\langle v\rangle+\langle x\rangle\big)|u|M_{\beta}(u)M_{\alpha}(y)|f^{\delta}-f|(x-y,v-u){\,\rm d}y{\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big|$
    $\leq\|\langle v\rangle^{2}M_{\beta}\|_{L^{1}}\|\langle x\rangle^{2}M_{\alpha}\|_{L^{1}}\Big|\int_{0}^{T}\int_{\mathbb{R}^{2d}}\big(\langle v\rangle+\langle x\rangle\big)(f^{\delta}_{t}-f_{t}){\,\rm d}x{\,\rm d}v{\,\rm d}t\Big|,$
    which vanishes as $\delta\to 0$, since $f\in L^{1}([0,T];L^{1}_{1,1}(\mathbb{R}^{2d}))$. By the dominated convergence theorem, we obtain the limit as $\delta\to 0$.
@@
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big|$
    $\leq$ ${}C(\alpha,\beta)\int_{s}^{t}\int_{\mathbb{R}^{4d}}\big(\langle v\rangle+\langle x\rangle+\mathcal{E}_{2,0}(f_{t})^{\frac{1}{2}}\big)$
    $\quad\times|u|M_{\beta}(u)M_{\alpha}(y)f(x-y,v-u){\,\rm d}y{\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
@@
    $\lim_{\beta\to 0}\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r=0.$
@@
    $\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}(v)\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $=-\int_{s}^{t}\int_{\mathbb{R}^{3d}}\frac{\nabla_{x}f^{\alpha,\beta}_{r}(v)}{f^{\alpha,\beta}_{r}(v)}\cdot uf^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r.$
@@
    $\Big|\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{\nabla_{x}f^{\alpha,\beta}_{r}(v)}{f^{\alpha,\beta}_{r}(v)}\cdot\int_{\mathbb{R}^{d}}uf^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big|$
    $\leq\Big(\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{|\nabla_{x}f^{\alpha,\beta}_{r}|^{2}}{f^{\alpha,\beta}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big)^{\frac{1}{2}}\times$
    $\quad\times\Big(\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{(\int_{\mathbb{R}^{d}}|u|f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u)^{2}}{\int_{\mathbb{R}^{d}}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u}{\,\rm d}x{\,\rm d}v{\,\rm d}r\Big)^{\frac{1}{2}}.$
@@
    $\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{(\int_{\mathbb{R}^{d}}|u|f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u)^{2}}{\int_{\mathbb{R}^{d}}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $\leq{}$ $\int_{s}^{t}\int_{\mathbb{R}^{3d}}\frac{(|u|f^{\alpha}_{r}(v-u))^{2}}{f^{\alpha}_{r}(v-u)}M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r$
    $\leq{}$ $\|f\|_{L^{1}([0,T]\times\mathbb{R}^{2d})}\int_{\mathbb{R}^{d}}{|v|^{2}M_{\beta}(v)}{\,\rm d}v$
    $\leq{}$ $CT\beta^{2}.$
@@
    $\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{|\nabla_{x}f^{\alpha,\beta}_{r}|^{2}}{f^{\alpha,\beta}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r$ $\leq\int_{s}^{t}\int_{\mathbb{R}^{2d}}\frac{|\nabla_{x}f^{\alpha}_{r}|^{2}}{f^{\alpha}_{r}}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
@@
    $|\int_{s}^{t}\int_{\mathbb{R}^{2d}}\log f^{\alpha,\beta}_{r}(v)\int_{\mathbb{R}^{d}}u\cdot\nabla_{x}f^{\alpha}_{r}(v-u)M_{\beta}(u){\,\rm d}u{\,\rm d}x{\,\rm d}v{\,\rm d}r|$
@@
$\leq{}$ $\int_{\mathbb{R}^{2d}}|f^{\alpha,\beta,\delta,\theta}_{t}\log f^{\alpha,\beta,\delta,\theta}_{t}-f_{t}^{\alpha,\beta,\theta}\log f_{t}^{\alpha,\beta,\theta}|{\,\rm d}x{\,\rm d}v$
$\leq{}$ $\int_{\mathbb{R}^{2d}}|f^{\alpha,\beta,\delta,\theta}_{t}-f_{t}^{\alpha,\beta,\theta}|\sup_{\delta}|\log f^{\alpha,\beta,\delta,\theta}_{t}|{\,\rm d}x{\,\rm d}v$
$\leq{}$ $C(\theta)\int_{\mathbb{R}^{2d}}\big(\langle v\rangle+\langle x\rangle\big)|f^{\alpha,\beta,\delta}_{t}-f_{t}^{\alpha,\beta}|{\,\rm d}x{\,\rm d}v\to 0$
@@
We left to show the $\alpha,\,\beta\to 0$ limits. Since $\mathcal{H}(f^{\alpha,\beta}_{t})$ is lower semicontinuous and $f^{\alpha,\beta}_{t}\to f_{t}$ in $L^{1}_{2,2}(\mathbb{R}^{2d})$ as $\alpha,\,\beta\to 0$, we have
@@
$\mathcal{H}\Big(\int_{\mathbb{R}^{2d}}f_{t}(x-y,v-u)M_{\alpha}(y)M_{\beta}(u){\,\rm d}y{\,\rm d}u\Big)\leq\mathcal{H}(f_{t}).$